OpenStudy (anonymous):

Verify the identity sin(pi/2-x)=sin(pi/2+x)

OpenStudy (alekos):

lhs=sin(pi/2 - x) = sin pi/2cosx - sinxcospi/2 = cosx rhs= sin(pi/2+x) = sinpi/2cosx + sinxcospi/2 = cosx sinpi/2=1 and cospi/2=0

OpenStudy (anonymous):

thank you!

OpenStudy (alekos):

no problem